# Performance and radiation damage mitigation strategy for silicon photomultipliers on LEO space missions

### v2

@@ 4 Evaluation of radiation fluxes for a LEO space mission
In order to evaluate the radiation damage in space on the sensors of the FPA, we use the SPENVIS (Space Environment Information System) [20] algorithm, a powerful software suite developed by the European Space Agency (ESA). In this framework, we built a project according to the specification of our mission’s orbit altitude of 550 km and an inclination of approximately $97.8^{\circ}$ between mid-2026 to mid-2029, which corresponds to an intense solar activity (see figure 13).
@@ 5 Dose estimates from radiation in space with Geant4
To estimate the dose that we expect on Terzina’s camera, we simulated the geometry of the telescope in Geant4. The simulated mechanics has small differences with respect to the final mechanics defined subsequently, but the few changes are irrelevant for this study. The simulated geometry, similar to the CAD in figure 2, includes:
@@
    a holder made of carbon fibre composed of a small baffle and a tower that protects the secondary mirror;
@@
    the focal plane composed by the sensors and the PCB connecting the tiles to the rest of the data acquisition system;
@@
We tested different geometrical configurations varying the thickness and height of the external baffle and the thickness of the small baffle, indicated in Fig. 2. In particular, we report the results for three configurations: the hereafter called baseline geometry, with an external baffle of thickness of 1 mm and maximum height of the transverse section of 730 mm and the small baffle of 1 mm thickness; the same external cut baffle and a small baffle 4 mm thick, called benchmark configuration; a better shielded configuration, called shielded, consisting of a cylindrical external baffle 2 mm thick and of height 730 mm without a transverse cut, and a small baffle 4 mm thick.
Protons and electrons have been injected into these geometries simulated in Geant4 to estimate the total radiation dose. Particles were simulated in small energy bins considering the full energy range of the flux calculated with SPENVIS $\frac{d\phi}{dE}$ integrated over a sphere of radius $R=50$ cm surrounding the considered volumes and for the duration of the mission $t_{\mathrm{EoL}}$ of 3 years. The total energy deposited in the sensitive volume of the FPA tiles and in the PCB, $E_{\mathrm{dep}}$ returned in MeV, has been calculated by Geant4. The total dose released in the corresponding volumes for all particles in the energy bin has been calculated summing all the differential doses:
@@
returned in J/kg or gray, which is the SI unit of ionizing radiation dose defined as the absorption of 1 J of radiation energy per kg the object. We account for the fact that each particle crosses the spherical volume from the two opposite sides of the sphere by dividing by 2 in eq. 5.1, assuming that volumes have a uniform density and that $M_{\mathrm{tot}}$ is their total mass (sensors or PCB), and $c_{\mathrm{MeVJ}}$ the conversion factor from MeV to Joule. These total doses for a 3-year mission, released in the simulated Terzina geometries by trapped electrons and protons, and solar protons are summarized in table 3. Considering these results, the shield power and the allowed weight for the mission, the benchmark configuration has been chosen for implementation, as the shielded one would not be affordable for weight limitations.
@@ 6 Radiation damage in silicon with proton irradiation
After the full characterisation of the sensors in section 2, we have irradiated them with protons and electrons in order to estimate the radiation damage. When an energetic charged particle hits a SiPM, it deposits energy through both ionizing and non-ionizing processes. Radiation damage of SiPMs has been extensively discussed in [40]. In silicon, bulk damage is due to Non-Ionizing Energy Loss (NIEL) and surface damage by Ionizing Energy Loss (IEL). Bulk damage is due to high energy particles (can be protons, neutrons, electrons, pions or photons) which can displace atoms out of their lattice creating defects as shown in figure 14. At energies beyond the keV scale, cluster defects can form when more atoms are displaced. Electrons beyond energies of 1 MeV typically can produce single defects, whereas heavier particles like protons produce also cluster defects. In the NIEL hypothesis, the radiation damage is proportional to the non-ionizing energy loss of the particles and this energy loss is proportional to the energy used to dislocate the lattice atoms (displacement energy). The defects in silicon crystals lead to an increase in leakage current due to the generation of electron-hole pairs from defects in the depletion region and trap-assisted tunneling. This leads to a temperature-dependent increase in DCR, which is more significant for protons than for electrons, which primarily induce ionisation rather than defects due to their lower mass and less efficient energy transfer to the crystal lattice, resulting in a minimal increase in DCR. A detail computation of the NIEL for protons and electrons in silicon can be done on Screened Relativistic (SR) Treatment for NIEL Dose Nuclear and Electronic Stopping Power Calculator [41]. Hence, we expect that damage caused by electrons induce a lower increase of the DCR when compared to protons at equal over-voltage, temperature, cell sizes, and dose.
@@
We performed the proton irradiation tests at the IFJ PAN in Krakow [42] with a 50 MeV proton beam (see the experimental setup in figure 15). The proton beam spot had a circular shape with a 35 mm diameter and homogeneity better than 5% with respect to the mean fluence. We measured the currents for $1\times 1$ mm2 SiPMs with $\mu$-cell size of 25 $\mu\text{m}$, 30 $\mu\text{m}$, 35 $\mu\text{m}$ and 50 $\mu\text{m}$ and 3$\times$3 mm2 SiPMs with a $\mu$-cell size of 40 $\mu\text{m}$ with and without resin coating. We measured the current of the SiPMs as a function of the bias voltage in the post-breakdown Geiger regime. The currents have been acquired just before the start of and during the irradiation process are shown as a function of the dose for the different devices in figure 16. The expected linear behaviour has been fitted as:
@@
where the fit parameters, $m$ and $I_{0}$, are reported in table 4 for $V_{\mathrm{bias}}=42$ V. The empirical model obtained from the fit will be used to predict the increase of the DCR as a function of dose in section 9.1. Variations of other SiPMs features, such as breakdown voltage and quenching resistance, have not been observed after irradiation.
@@ 7 Thermal effect
To characterize the behaviour of the sensors as a function of temperature, we measured IV curves at different temperatures and extrapolated the DCR based on these curves. DCR being proportional to the post-breakdown current at a given over-voltage, we can derive the DCR from the current measurements. The results are used to select the sensor and set the operating temperature. Tests were performed both before and after irradiation. The temperature dependency of the IV curve measurements is modelled with the function, with parameters $I_{0}$ and $\lambda$:
$I(T)=I_{0}\cdot e^{\lambda T},$ (7.1)
The fit results are reported in table 5 and the measurements and the fit curves are shown in figure 17. It can be noted that the obtained parameter $I_{0}$ corresponds to the current at $T(^{\circ}\mathrm{C})=0$ and the rate constant $\lambda$ related to the exponential temperature constant $\tau=1/\lambda$ provides the temperature correction factor used to obtain the currents at the desired temperature throughout this work. We can estimate the change in current per degree in order to get the temperature correction that we used in all the measurements. For example, we can estimate that, if the current value is doubled, the temperature will be obtained as $2I_{0}=I_{0}\cdot e^{\lambda T_{I_{2}}}$ as: $T_{I_{2}}=\frac{\ln{2}}{\lambda}$. The current is doubled every 10 degrees before irradiation and every 15 degrees after irradiation (see last column in table 5). As shown in figure 17, the increase in DCR with temperature is slightly slower than in the presence of radiation.
Figure 17: Measured currents at $V_{\mathrm{OV}}=8.5$ V as a function of temperature. The lines represent the fit results described by equation 7.1.
@@ 8 Electrons irradiation measurements and simulations
In order to measure the radiation damage effect on SiPMs, namely the increase of DCR due to electrons, and predict this effect during the NUSES mission, we used a strontium 90 (${}^{90}Sr$) $\beta$ emitting source with activity of 16.75 MBq. Two bare and coated NUV-HD-MT of $3\times 3$ mm2 SiPMs with 40 $\mu\text{m}$-cell size and $8.52\,\text{mm}^{2}$ sensitive area were exposed to the source. All measurements are done in an over-voltage interval from 0 to 18 V. Considering the energy spectrum of primary and secondary electrons from ${}^{90}Sr$ decays [43], with this source we can study radiation damage from electrons up to a maximum energy of 2.28 MeV. As done for protons, we measured the current increase, but in this case we moved the source at different distances from the SiPM using the experimental setup shown in figure 18. We also built up a full simulation in Geant4 of this setup to compare with the measurements. Figure 19 depicts the measured current as a function of time at different over-voltages and changing the distance of the source from the SiPM as indicated on the top axis.
@@
Figure 19: The current as a function of time for three over-voltage values after temperature correction (using equation 7.1), for bare SiPMs (left) and coated ones (right) . Each indicated time range with different colours and on the top horizontal axis corresponds to the indicated distance between the source and SiPM starting from no source ($\infty$). The zoomed plot in the insert on the right panel for the coated SiPM shows one of the resulting Cherenkov jumps due to moving the source closer to the SiPM.
@@
Figure 20 shows the same current as a function of over-voltage (IV curves) for the same sensors, with bare SiPM on the left panel and coated SiPM on the right panel. The time of irradiation for each IV curve is indicated in the label. First, we put the source at a 12 mm distance to the SiPM (lower dashed curves). It can be seen that the different times of exposure to the source introduce a small difference due to the electron radiation damage for different exposure times. Then we move the source to 8 mm distance to the SiPM and we again see that for different exposure times the effect on the IV curve is small between, but on the right panel for the coated SiPM there is a larger increase as we moved the source closer to the SiPM. This indicates that there is an additional contribution to the radiation damage of electrons which is due to the Cherenkov emission of electrons in the coating resin. As expected, this effect is not observed on the left panel for the bare SiPM.
@@
Figure 20: The IV curve for the same bare SiPM (left) and coated SiPM (right), measured 4 different times. The lower 2 curves are measured with the source at a 12 mm distance to the SiPM ($\Delta r=12$ mm), while the other 2 were taken with the source positioned 8 mm away from the SiPM ($\Delta r=8$ mm). The corresponding dose is included in the plot by using the results from the simulation and equation 8.1.
@@
The indicated doses in figure 20 are calculated as follows. The simulation provides us with the precise energy deposition ($dE/dx$) of electrons resulting from ionizing and non-ionizing processes as they pass through each layer of the SiPM. By dividing the silicon volume into 8 layers and running the simulation many times, we obtained the mean value of the total energy deposit in the volume of the SiPM, denoted as $dE/dx|_{\text{mean}}$. We can then compare the $dE/dx|_{\text{mean}}$ or dose in each layer to analyse where most of the OCT and DCR contributed. Using the following formula, we calculate the dose ($\mathcal{D}$) in each layer:
@@
Here, $e$, $\mathcal{A}$, and $m_{\text{V}}$ represent the charge of an electron, the activity of the source, and the mass of the silicon layer, respectively. It is worth noting that each simulation iteration corresponds to a single decay event of ${}^{90}Sr$. Considering the source’s activity, this equation provides the equivalent of a $dE/dx|_{\text{mean}}$ per one second, and thus, we derive the dose rate in Gy/hr. Figure 21 shows the final dose from the simulation in each silicon layer of the SiPM.
@@ 9 DCR prediction during mission
The DCR represents the baseline count rate in the absence of incident light. In order to avoid that the camera is constantly triggering on DCR induced events, the threshold of the trigger has to be increased with increasing DCR. This impacts the lowest achievable energy of the detected primary cosmic rays. With all the measurements that have been done for thermal response, electrons and protons campaigns, we can now calculate the DCR considering the overall effects through the following formula:
@@
Here, the $V_{\textrm{ov}}$, $\mathcal{S}$, Gain, $T_{\mathrm{op}}$, $\mathcal{D}$, and $q$ stand for over-voltage ($V_{\textrm{ov}}=10$ V), effective sensitive area, gain of the SiPM in use, average operation temperature during the mission ($T_{\mathrm{op}}=0^{\circ}$), dose released by either protons or electrons, and the electron charge, respectively.
The $I^{\ast}(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)=\frac{I(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)}{s^{\ast}}$ corresponds to the measured current $I(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)$ divided by $s^{\ast}$, namely scaling for the ratio of areas to take into account the different SiPM sizes. The exponential term provides the measured current at desired temperature following the model in equation 7.1. Combining the DCR results with the dose-per-time response from the simulation yields the results in figure 22. As can be seen from the graph, the damage caused by electrons is very low when compared to protons at equal over-voltage, temperature, cell sizes, and dose as expected. Specifically, 2 Gy of electrons corresponds to a DCR of approximately 0.15 MHz, while 2 Gy from protons results in a DCR of around 50 MHz, starting from the same DCR at 0 Gy.
Figure 22: Inferred DCR per pixel of Terzina SiPMs with sensitive area of $\sim 6.58\leavevmode\nobreak\ \text{mm}^{2}$ from IV measurements during the irradiation as a function of dose and mission timeline. The errors are due to the instrument scale precision that changes with range values. The change of scale is clearly visible.
For this reason in the next section only the radiation damage produced by protons will be taken into account in the thermal annealing strategy done to recover the SiPM response.

### v3

@@ 4 Evaluation of radiation fluxes for a LEO space mission
In order to evaluate the radiation damage in space on the sensors of the FPA, we used the SPENVIS (Space Environment Information System) [20] algorithm, a powerful software suite developed by the European Space Agency (ESA). In this framework, we built a project according to the specification of our mission’s orbit altitude of 550 km and an inclination of approximately $97.8^{\circ}$ between mid-2026 to mid-2029, which corresponds to an intense solar activity (see figure 13).
@@ 5 Dose estimates from radiation in space with Geant4
To estimate the dose that we expect on Terzina’s camera, we simulated the geometry of the telescope in Geant4. The simulated mechanics has small differences with respect to the final one defined subsequently, but these few changes are irrelevant for this study. The simulated geometry, similar to the CAD in figure 2, includes:
@@
    a holder made of carbon fibre, composed of a small baffle and a tower that protects the secondary mirror;
@@
    the focal plane composed of the sensors and the PCB connecting the tiles to the rest of the data acquisition system;
@@
We tested different geometrical configurations varying the thickness and height of the external baffle and the thickness of the small baffle, indicated in figure 2. In particular, we report the results for three configurations: the hereafter called baseline geometry, with an external baffle of thickness of 1 mm and maximum height of the transverse section of 730 mm and the small baffle of 1 mm thickness; the same external cut baffle and a small baffle 4 mm thick, called benchmark configuration; a better shielded configuration, called shielded, consisting of a cylindrical external baffle 2 mm thick and of height 730 mm without a transverse cut, and a small baffle 4 mm thick.
Protons and electrons have been injected into these geometries, simulated in Geant4, to estimate the total radiation dose. Particles were simulated in small energy bins considering the full energy range of the flux calculated with SPENVIS $\frac{d\phi}{dE}$ integrated over a sphere of radius $R=50$ cm surrounding the considered volumes and for the duration of the mission $t_{\mathrm{EoL}}$ of 3 years. The total energy deposited in the sensitive volume of the FPA tiles and in the PCB, $E_{\mathrm{dep}}$ returned in MeV, has been calculated by Geant4. The total dose released in the corresponding volumes for all particles in the energy bin has been calculated by summing all the differential doses:
@@
returned in J/kg or Gray, which is the SI unit of ionizing radiation dose defined as the absorption of 1 J of radiation energy per kg of the object. We account for the fact that each particle crosses the spherical volume from the two opposite sides of the sphere by dividing by 2 in eq. 5.1, assuming that volumes have a uniform density and that $M_{\mathrm{tot}}$ is their total mass (sensors or PCB), and $c_{\mathrm{MeVJ}}$ the conversion factor from MeV to Joule. These total doses for a 3-year mission, released in the simulated Terzina geometries by trapped electrons and protons, and solar protons are summarized in table 3. Considering these results, the shield power and the allowed weight for the mission, the benchmark configuration has been chosen for implementation, as the shielded one would not be affordable for weight limitations.
@@ 6 Radiation damage in silicon with proton irradiation
After the full characterisation of the sensors in section 2, we have irradiated them with protons and electrons in order to estimate the radiation damage. When an energetic charged particle hits a SiPM, it deposits energy through both ionizing and non-ionizing processes. Radiation damage of SiPMs has been extensively discussed in [40]. In silicon, bulk damage is due to Non-Ionizing Energy Loss (NIEL) and surface damage by Ionizing Energy Loss (IEL). Bulk damage is due to high-energy particles (can be protons, neutrons, electrons, pions or photons), which can displace atoms out of their lattice creating defects as shown in figure 14. At energies beyond the keV scale, cluster defects can form when more atoms are displaced. Electrons beyond energies of 1 MeV typically can produce single defects, whereas heavier particles like protons also produce cluster defects. In the NIEL hypothesis, the radiation damage is proportional to the non-ionizing energy loss of the particles and this energy loss is proportional to the energy used to dislocate the lattice atoms (displacement energy). The defects in silicon crystals lead to an increase in leakage current due to the generation of electron-hole pairs from defects in the depletion region and trap-assisted tunnelling. This leads to a temperature-dependent increase in DCR, which is more significant for protons than for electrons, which primarily induce ionisation rather than defects due to their lower mass and less efficient energy transfer to the crystal lattice, resulting in a minimal increase in DCR. A detailed computation of the NIEL for protons and electrons in silicon can be done with the Screened Relativistic (SR) Treatment for NIEL Dose Nuclear and Electronic Stopping Power Calculator [41]. Hence, we expect that the damage caused by electrons induce a lower increase in DCR when compared to protons at equal over-voltage, temperature, cell sizes, and dose.
@@
We performed the proton irradiation tests at the IFJ PAN in Krakow [42] with a 50 MeV proton beam (see the experimental setup in figure 15). The proton beam spot had a circular shape with a 35 mm diameter and homogeneity better than 5% with respect to the mean fluence. We measured the currents for $1\times 1$ mm2 SiPMs with $\mu$-cell size of 25 $\mu\text{m}$, 30 $\mu\text{m}$, 35 $\mu\text{m}$ and 50 $\mu\text{m}$ and 3$\times$3 mm2 SiPMs with a $\mu$-cell size of 40 $\mu\text{m}$ with and without resin coating. We measured the current of the SiPMs as a function of the bias voltage in the post-breakdown Geiger regime. The currents acquired just before the start of and during the irradiation process are shown as a function of the dose for the different devices in figure 16. The expected linear behaviour has been fitted as:
@@
where the fit parameters, $m$ and $I_{0}$, are reported in table 4 for $V_{\mathrm{bias}}=42$ V. The empirical model obtained from the fit will be used to predict the increase of the DCR as a function of dose in section 9.1. Variations of other SiPM features, such as breakdown voltage and quenching resistance, have not been observed after irradiation.
@@ 7 Thermal effect
The optimal operation temperature for Terzina is considered to be -20C and the thermal control will avoid temperatures larger than 5C. In this work, we performed the sensor characterisation at room temperature (see section 3). Nonetheless, we performed some useful IV measurements in a climate chamber to characterise the behaviour of the sensors as a function of their temperature. These allowed us to extrapolate the DCR behaviour at other temperatures than room temperature. Being the DCR proportional to the post-breakdown current at a given over-voltage, we can derive the DCR from the current measurements. The results are used to select the sensors and understand if it was necessary any action to correct the operation temperature. Tests were performed both before and after irradiation. The temperature dependency of the IV curve measurements is modelled with the function, with parameters $I_{0}$ and $\lambda$:
$I(T)=I_{0}\cdot e^{\lambda T}\,.$ (7.1)
The fit results are reported in table 5 and the measurements and the fit curves are shown in figure 17. It can be noted that the obtained parameter, $I_{0}$, corresponds to the current at $T(^{\circ}\mathrm{C})=0$ and the rate constant $\lambda$ related to the exponential temperature constant $\tau=1/\lambda$ provides the temperature correction factor used to obtain the currents at the desired temperature throughout this work. We can estimate the change in current per degree in order to get the temperature correction that we used in all the measurements. For example, we can estimate that, if the current value is doubled, the temperature will be obtained as $2I_{0}=I_{0}\cdot e^{\lambda T_{I_{2}}}$ as: $T_{I_{2}}=\frac{\ln{2}}{\lambda}$. The current is doubled every 10 degrees before irradiation and every 15 degrees after irradiation (see last column in table 5). As shown in figure 17, the increase in DCR with temperature is slightly slower than in the presence of radiation.
Figure 17: Measured currents (symbols with statistical errors) at $V_{\mathrm{OV}}=8.5$ V as a function of temperature. The lines represent the fit results described by equation 7.1.
@@ 8 Electrons irradiation measurements and simulations
In order to measure the radiation damage effect on SiPMs, namely the increase of DCR due to electrons, and predict this effect during the NUSES mission, we used a strontium 90 (${}^{90}Sr$) $\beta$ emitting source with activity of 16.75 MBq. Two bare and coated (resin 0.75 mm thick) NUV-HD-MT of $3\times 3$ mm2 SiPMs with 40 $\mu\text{m}$-cell size and $8.52\,\text{mm}^{2}$ sensitive area were exposed to the source. All measurements are done in an over-voltage interval from 0 to 18 V. Considering the energy spectrum of primary and secondary electrons from ${}^{90}Sr$ decays [43], with this source, we can study radiation damage from electrons up to a maximum energy of 2.28 MeV. As done for protons, we measured the current increase, but in this case, we moved the source at different distances from the SiPM using the experimental setup shown in figure 18. We also built up a full simulation in Geant4 of this setup to compare with the measurements. Figure 19 depicts the measured current as a function of time at different over-voltages and changing the distance of the source from the SiPM as indicated on the top axis.
@@
Figure 19: The current as a function of time for three over-voltage values after temperature correction (using equation 7.1), for bare SiPMs (left) and coated ones (right) . Each indicated time range, highlighted with different colours and on the top horizontal axis, corresponds to the indicated distance between the source and SiPM starting from no source ($\infty$). The zoomed plot in the insert on the right panel for the coated SiPM shows one of the resulting Cherenkov jumps due to moving the source closer to the SiPM.
@@
Figure 20 shows the same current as a function of over-voltage (IV curves) for the same sensors, with bare SiPM on the left panel and coated SiPM on the right panel. The time of irradiation for each IV curve is indicated on the label. First, we put the source at a 12 mm distance from the SiPM (lower dashed curves). It can be seen that the different times of exposure to the source introduce a small difference due to the electron radiation damage for different exposure times. Then we moved the source to 8 mm distance to the SiPM and we again see that for different exposure times the effect on the IV curve is small between, but on the right panel for the coated SiPM there is a larger increase as we moved the source closer to the SiPM. This indicates that, other than radiation damage, there is an additional contribution to the current increase, which is due to the Cherenkov emission of electrons in the coating resin. As expected, this effect is not observed in the left panel for the bare SiPM.
@@
Figure 20: The IV curve for the same bare SiPM (left) and coated SiPM (right), measured at different times. The lower two curves are measured with the source at a 12 mm distance from the SiPM ($\Delta r=12$ mm), while the other two were taken with the source positioned 8 mm away from the SiPM ($\Delta r=8$ mm). The corresponding dose is included in the plot by using the results from the simulation and equation 8.1.
@@
The indicated doses in figure 20 are calculated as follows. The simulation provides us with the precise energy deposition ($dE/dx$) of electrons resulting from ionizing and non-ionizing processes as they pass through each layer of the SiPM. By dividing the silicon volume into 8 layers (with the thickness of 60 $\mu$m for each layer) and running the simulation many times, we obtained the mean value of the total energy deposit in the volume of the SiPM, denoted as $dE/dx|_{\text{mean}}$. We can then compare the $dE/dx|_{\text{mean}}$ or dose in each layer to analyse where most of the OCT and DCR contributed. Using the following formula, we calculate the dose ($\mathcal{D}$) in each layer:
@@
Here, $e$, $\mathcal{A}$, and $m_{\text{V}}$ represent the charge of an electron, the activity of the source, and the mass of the silicon layer, respectively. It is worth noting that each simulation iteration corresponds to a single decay event of ${}^{90}Sr$. Considering the source’s activity, this equation provides the equivalent of a $dE/dx|_{\text{mean}}$ per one second, and thus, we derive the dose rate in Gy/hr. Figure 21 shows the final dose from the simulation in each silicon layer of the SiPM. The upper layers absorb the highest dose and thus experience the most radiation damage. Additionally, comparing the bare and coated SiPM configurations reveals that the 0.75 mm-thick resin on top of the SiPM reduces the dose received by the layers of the SiPM.
@@ 9 DCR prediction during mission
The DCR represents the baseline count rate in the absence of incident light. In order to avoid that the camera is constantly triggering on DCR induced events, the threshold of the trigger has to be increased with increasing DCR. This impacts the lowest achievable energy of the detected primary cosmic rays. With all the measurements that have been done for thermal response, electrons and protons campaigns, we can now calculate the DCR at a certain operation temperature (e.g. at $T_{\mathrm{op}}=0^{\circ}$) considering the overall effects through the following formula:
@@
Here, the $V_{\textrm{ov}}$, $\mathcal{S}$, Gain, $T_{\mathrm{op}}$, $\mathcal{D}$, and $q$ stand for over-voltage ($V_{\textrm{ov}}=10$ V), effective sensitive area, gain of the SiPM in use, average operation temperature during the mission, dose released by either protons or electrons, and the electron charge, respectively.
The $I^{\ast}(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)=\frac{I(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)}{s^{\ast}}$ corresponds to the measured current $I(\textrm{V${}_{\textrm{ov}}$},\text{$\mathcal{D}$},T)$ divided by $s^{\ast}$, namely scaling for the ratio of areas to take into account the different SiPM sizes. The exponential term provides the measured current at desired temperature following the model in equation 7.1. Combining the DCR results with the dose-per-time response from the simulation yields the results in figure 22. As can be seen from the graph, the damage caused by electrons is very low when compared to protons at equal over-voltage, temperature, cell sizes, and dose, as expected. Specifically, 2 Gy of electrons corresponds to a DCR of approximately 0.15 MHz, while 2 Gy from protons results in a DCR of around 50 MHz, starting from the same DCR at 0 Gy.
Figure 22: Inferred DCR per pixel of Terzina SiPMs with sensitive area of $\sim 6.58\leavevmode\nobreak\ \text{mm}^{2}$ from IV measurements during the irradiation as a function of dose and mission timeline. The errors are due to the instrument scale precision, which changes with range as can be seen in the plot.
For this reason, in the next section, only the radiation damage produced by protons will be taken into account in the thermal annealing strategy done to recover the SiPM response.